OpenStudy (anonymous):

What is a possible value for the missing term of the geometric sequence? 50, ___, 450, ...

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf 50 \times X=450 \div X }\) where x is the common ratio then, what each of the sides (in my original equation) is equal to, is the missing term. or so called, "geometric mean"

OpenStudy (texaschic101):

you can make a common difference of 3

OpenStudy (texaschic101):

oops....common ratio of 3

OpenStudy (anonymous):

How did you get the common ratio? 450/50 = 9

OpenStudy (anonymous):

a,b,c are in G.P. THEN\[\frac{ b }{ a }=\frac{ c }{b },or~b ^{2}=ac\]

OpenStudy (anonymous):

\[\{ if x is the \middle term ,then x ^{2}=50*450}\] \[x=\pm \sqrt{50*450}=\pm 50*3=\pm 150\] if x=150 ,c.r.=150/50=3 terms are 50,150,450,1350,.... if x=-150,c.r.=-150/50=-3 terms are 50,-150,450,-1350,... terms are alternately positive and negative.